OpenStudy (anonymous):

Inflation is causing a rise of approximately 2% per year in the cost of living. A bag of milk costs $3.95 now. Estimate the cost five years from now.

OpenStudy (anonymous):

1 year .02(3.95)+3.95 = 4.029 2 years .02(4.029)+4.029=4.10958 3 years .02(4.10958)+4.10958=4.1917716 4 years .02(4.1917716)+4.1917716=4.275607032 5 years .02(4.275607032)+4.275607032=4.36111917264 After 5 years it would become $4.36

OpenStudy (mathstudent55):

Now: $3.95 = $3.95 In 1 year: $3.95 * 1.02 = $3.95 * 1.02^1 In 2 years: ($3.95 * 1.02) * 1.02 = $3.95 * 1.02^2 In 3 years: (3.95 * 1.02 * 1.02) * 1.02 = $3.95 * 1.02^3 In 4 years: ($3.95 * 1.02 * 1.02 * 1.02) * 1.02 = $3.95 * 1.02^4 In n years: = $3.95 * 1.02^n

OpenStudy (mathstudent55):

In your case, for 5 years, $3.95 * 1.02^5 = $4.36